\begin{thebibliography}{40}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Barnard(2003)]{Bar03}
Alexander~Graham Barnard.
\newblock \emph{The singular theta correspondence, {L}orentzian lattices and
  {B}orcherds-{K}ac-{M}oody algebras}.
\newblock ProQuest LLC, Ann Arbor, MI, 2003.
\newblock Thesis (Ph.D.)--University of California, Berkeley.

\bibitem[Borcherds(1998)]{Bor98}
Richard Borcherds.
\newblock Automorphic forms with singularities on {G}rassmannians.
\newblock \emph{Invent. Math.}, 132\penalty0 (3):\penalty0 491--562, 1998.

\bibitem[Borcherds(1992)]{Bor92}
Richard~E. Borcherds.
\newblock Monstrous moonshine and monstrous {L}ie superalgebras.
\newblock \emph{Invent. Math.}, 109\penalty0 (2):\penalty0 405--444, 1992.

\bibitem[Borcherds(1995)]{Bor95}
Richard~E. Borcherds.
\newblock Automorphic forms on {${\rm O}_{s+2,2}({\bf R})$} and infinite
  products.
\newblock \emph{Invent. Math.}, 120\penalty0 (1):\penalty0 161--213, 1995.

\bibitem[Borcherds(1996)]{Bor96}
Richard~E. Borcherds.
\newblock The moduli space of {E}nriques surfaces and the fake {M}onster {L}ie
  superalgebra.
\newblock \emph{Topology}, 35\penalty0 (3):\penalty0 699--710, 1996.

\bibitem[Borcherds(2000)]{Bor00}
Richard~E. Borcherds.
\newblock Reflection groups of {L}orentzian lattices.
\newblock \emph{Duke Math. J.}, 104\penalty0 (2):\penalty0 319--366, 2000.

\bibitem[Borcherds et~al.(1998)Borcherds, Katzarkov, Pantev, and
  Shepherd-Barron]{BKP98}
Richard~E. Borcherds, Ludmil Katzarkov, Tony Pantev, and N.~I. Shepherd-Barron.
\newblock Families of {$K3$} surfaces.
\newblock \emph{J. Algebraic Geom.}, 7\penalty0 (1):\penalty0 183--193, 1998.

\bibitem[Bruinier(2002)]{Bru02}
Jan~H. Bruinier.
\newblock \emph{Borcherds products on {O}(2, {$l$}) and {C}hern classes of
  {H}eegner divisors}, volume 1780 of \emph{Lecture Notes in Mathematics}.
\newblock Springer-Verlag, Berlin, 2002.

\bibitem[Bruinier(2014)]{Bru14}
Jan~Hendrik Bruinier.
\newblock On the converse theorem for {B}orcherds products.
\newblock \emph{J. Algebra}, 397:\penalty0 315--342, 2014.

\bibitem[Conway and Sloane(1999)]{CS99}
J.~H. Conway and N.~J.~A. Sloane.
\newblock \emph{Sphere packings, lattices and groups}, volume 290 of
  \emph{Grundlehren der mathematischen Wissenschaften [Fundamental Principles
  of Mathematical Sciences]}.
\newblock Springer-Verlag, New York, third edition, 1999.
\newblock With additional contributions by E. Bannai, R. E. Borcherds, J.
  Leech, S. P. Norton, A. M. Odlyzko, R. A. Parker, L. Queen and B. B. Venkov.

\bibitem[Dittmann(2019)]{Dit18}
Moritz Dittmann.
\newblock Reflective automorphic forms on lattices of squarefree level.
\newblock \emph{Trans. Amer. Math. Soc.}, 372\penalty0 (2):\penalty0
  1333--1362, 2019.

\bibitem[Eichler and Zagier(1985)]{EZ85}
Martin Eichler and Don Zagier.
\newblock \emph{The theory of {J}acobi forms}, volume~55 of \emph{Progress in
  Mathematics}.
\newblock Birkh\"{a}user Boston, Inc., Boston, MA, 1985.

\bibitem[Gritsenko and Hulek(2014)]{GH14}
V.~Gritsenko and K.~Hulek.
\newblock Uniruledness of orthogonal modular varieties.
\newblock \emph{J. Algebraic Geom.}, 23\penalty0 (4):\penalty0 711--725, 2014.

\bibitem[Gritsenko et~al.(2009)Gritsenko, Hulek, and Sankaran]{GHS09}
V.~Gritsenko, K.~Hulek, and G.~K. Sankaran.
\newblock Abelianisation of orthogonal groups and the fundamental group of
  modular varieties.
\newblock \emph{J. Algebra}, 322\penalty0 (2):\penalty0 463--478, 2009.

\bibitem[Gritsenko et~al.(2013)Gritsenko, Hulek, and Sankaran]{GHS13}
V.~Gritsenko, K.~Hulek, and G.~K. Sankaran.
\newblock Moduli of {K}3 surfaces and irreducible symplectic manifolds.
\newblock In \emph{Handbook of moduli. {V}ol. {I}}, volume~24 of \emph{Adv.
  Lect. Math. (ALM)}, pages 459--526. Int. Press, Somerville, MA, 2013.

\bibitem[Gritsenko(2018)]{Gri18}
V.~A. Gritsenko.
\newblock Reflective modular forms and their applications.
\newblock \emph{Uspekhi Mat. Nauk}, 73\penalty0 (5(443)):\penalty0 53--122,
  2018.

\bibitem[Gritsenko and Nikulin(2002)]{GN02}
V.~A. Gritsenko and V.~V. Nikulin.
\newblock On the classification of {L}orentzian {K}ac-{M}oody algebras.
\newblock \emph{Uspekhi Mat. Nauk}, 57\penalty0 (5(347)):\penalty0 79--138,
  2002.

\bibitem[Gritsenko et~al.(2007)Gritsenko, Hulek, and Sankaran]{GHS07}
V.~A. Gritsenko, K.~Hulek, and G.~K. Sankaran.
\newblock The {K}odaira dimension of the moduli of {$K3$} surfaces.
\newblock \emph{Invent. Math.}, 169\penalty0 (3):\penalty0 519--567, 2007.

\bibitem[Gritsenko and Nikulin(1996)]{GN96}
Valeri~A. Gritsenko and Viacheslav~V. Nikulin.
\newblock {$K3$} surfaces, {L}orentzian {K}ac-{M}oody algebras and mirror
  symmetry.
\newblock \emph{Math. Res. Lett.}, 3\penalty0 (2):\penalty0 211--229, 1996.

\bibitem[Gritsenko and Nikulin(1998{\natexlab{a}})]{GN98a}
Valeri~A. Gritsenko and Viacheslav~V. Nikulin.
\newblock Automorphic forms and {L}orentzian {K}ac-{M}oody algebras. {I}.
\newblock \emph{Internat. J. Math.}, 9\penalty0 (2):\penalty0 153--199,
  1998{\natexlab{a}}.

\bibitem[Gritsenko and Nikulin(1998{\natexlab{b}})]{GN98b}
Valeri~A. Gritsenko and Viacheslav~V. Nikulin.
\newblock Automorphic forms and {L}orentzian {K}ac-{M}oody algebras. {II}.
\newblock \emph{Internat. J. Math.}, 9\penalty0 (2):\penalty0 201--275,
  1998{\natexlab{b}}.

\bibitem[Gritsenko and Nikulin(2000)]{GN00}
Valeri~A. Gritsenko and Viacheslav~V. Nikulin.
\newblock The arithmetic mirror symmetry and {C}alabi-{Y}au manifolds.
\newblock \emph{Comm. Math. Phys.}, 210\penalty0 (1):\penalty0 1--11, 2000.

\bibitem[Gritsenko and Nikulin(2018)]{GN18}
Valery Gritsenko and Viacheslav~V. Nikulin.
\newblock Lorentzian {K}ac-{M}oody algebras with {W}eyl groups of
  2-reflections.
\newblock \emph{Proc. Lond. Math. Soc. (3)}, 116\penalty0 (3):\penalty0
  485--533, 2018.

\bibitem[Looijenga(2003)]{Loo03}
Eduard Looijenga.
\newblock Compactifications defined by arrangements. {II}. {L}ocally symmetric
  varieties of type {IV}.
\newblock \emph{Duke Math. J.}, 119\penalty0 (3):\penalty0 527--588, 2003.

\bibitem[Ma(2017)]{Ma17}
Shouhei Ma.
\newblock Finiteness of 2-reflective lattices of signature {$(2,n)$}.
\newblock \emph{Amer. J. Math.}, 139\penalty0 (2):\penalty0 513--524, 2017.

\bibitem[Ma(2018)]{Ma18}
Shouhei Ma.
\newblock On the {K}odaira dimension of orthogonal modular varieties.
\newblock \emph{Invent. Math.}, 212\penalty0 (3):\penalty0 859--911, 2018.

\bibitem[Nikulin(1979)]{Nik80}
V.~V. Nikulin.
\newblock Integer symmetric bilinear forms and some of their geometric
  applications.
\newblock \emph{Izv. Akad. Nauk SSSR Ser. Mat.}, 43\penalty0 (1):\penalty0
  111--177, 238, 1979.

\bibitem[Nikulin(1980)]{Nik80b}
V.~V. Nikulin.
\newblock On the arithmetic groups generated by reflections in
  {L}oba\v{c}evski\u{\i} spaces.
\newblock \emph{Izv. Akad. Nauk SSSR Ser. Mat.}, 44\penalty0 (3):\penalty0
  637--669, 719--720, 1980.

\bibitem[Nikulin(1981)]{Nik81}
V.~V. Nikulin.
\newblock Quotient-groups of groups of automorphisms of hyperbolic forms by
  subgroups generated by {$2$}-reflections. {A}lgebro-geometric applications.
\newblock In \emph{Current problems in mathematics, {V}ol. 18}, pages 3--114.
  Akad. Nauk SSSR, Vsesoyuz. Inst. Nauchn. i Tekhn. Informatsii, Moscow, 1981.

\bibitem[Nikulin(1984)]{Nik84}
V.~V. Nikulin.
\newblock {$K3$} surfaces with a finite group of automorphisms and a {P}icard
  group of rank three.
\newblock volume 165, pages 119--142. 1984.
\newblock Algebraic geometry and its applications.

\bibitem[Nikulin(1996)]{Nik96}
V.~V. Nikulin.
\newblock Reflection groups in {L}obachevski\u{\i} spaces and an identity for
  the denominator of {L}orentzian {K}ac-{M}oody algebras.
\newblock \emph{Izv. Ross. Akad. Nauk Ser. Mat.}, 60\penalty0 (2):\penalty0
  73--106, 1996.

\bibitem[Pjatecki\u{\i}-\v{S}apiro and \v{S}afarevi\v{c}(1971)]{PS71}
I.~I. Pjatecki\u{\i}-\v{S}apiro and I.~R. \v{S}afarevi\v{c}.
\newblock Torelli's theorem for algebraic surfaces of type {${\rm K}3$}.
\newblock \emph{Izv. Akad. Nauk SSSR Ser. Mat.}, 35:\penalty0 530--572, 1971.

\bibitem[Scheithauer(2006)]{Sch06}
Nils~R. Scheithauer.
\newblock On the classification of automorphic products and generalized
  {K}ac-{M}oody algebras.
\newblock \emph{Invent. Math.}, 164\penalty0 (3):\penalty0 641--678, 2006.

\bibitem[Scheithauer(2017)]{Sch17}
Nils~R. Scheithauer.
\newblock Automorphic products of singular weight.
\newblock \emph{Compos. Math.}, 153\penalty0 (9):\penalty0 1855--1892, 2017.

\bibitem[Vinberg(1984)]{Vin84}
\`E.~B. Vinberg.
\newblock Discrete reflection groups in {L}obachevsky spaces.
\newblock In \emph{Proceedings of the {I}nternational {C}ongress of
  {M}athematicians, {V}ol. 1, 2 ({W}arsaw, 1983)}, pages 593--601. PWN, Warsaw,
  1984.

\bibitem[Vinberg(2007)]{Vin07}
\`E.~B. Vinberg.
\newblock Classification of 2-reflective hyperbolic lattices of rank 4.
\newblock \emph{Tr. Mosk. Mat. Obs.}, 68:\penalty0 44--76, 2007.

\bibitem[Wang(2019)]{Wan19}
Haowu Wang.
\newblock The classification of 2-reflective modular forms.
\newblock to appear in J. Eur. Math. Soc. (JEMS)., 2019.
\newblock URL \url{arXiv:1906.10459}.

\bibitem[Wang(2021{\natexlab{a}})]{Wan18}
Haowu Wang.
\newblock Reflective modular forms: a {J}acobi forms approach.
\newblock \emph{Int. Math. Res. Not. IMRN}, \penalty0 (3):\penalty0 2081--2107,
  2021{\natexlab{a}}.

\bibitem[Wang(2021{\natexlab{b}})]{Wan21}
Haowu Wang.
\newblock The classification of free algebras of orthogonal modular forms.
\newblock \emph{Compos. Math.}, 157\penalty0 (9):\penalty0 2026--2045,
  2021{\natexlab{b}}.

\bibitem[Wang(2022)]{Wan22}
Haowu Wang.
\newblock Reflective modular forms on lattices of prime level.
\newblock \emph{Trans. Amer. Math. Soc.}, 375\penalty0 (5):\penalty0
  3451--3468, 2022.

\end{thebibliography}

\documentclass[11pt]{amsart}

\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{amscd}
\usepackage{amsthm}
\usepackage{framed}
\usepackage{fullpage}
\usepackage{graphicx}
\usepackage{latexsym}
\usepackage[numbers]{natbib}  

\usepackage{hyperref}
\hypersetup{colorlinks,linkcolor={red},citecolor={blue},urlcolor={blue}}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{corollary}[theorem]{Corollary}

\theoremstyle{definition}

\numberwithin{equation}{section}

\newcommand{\CC}{\mathbb C}

\newcommand{\cD}{\mathcal D}
\newcommand{\cA}{\mathcal A}

\newcommand{\PP}{\mathbb P}

\newcommand{\ZZ}{\mathbb Z}

\newcommand{\Orth}{\mathop{\null\mathrm {O}}\nolimits}

\newcommand{\rank}{\mathop{\mathrm {rk}}\nolimits}
\newcommand{\latt}[1]{{\langle{#1}\rangle}}

\def\det{\operatorname{det}}

\begin{document}

\title[$2$-reflective lattices of signature $(n,2)$ with $n\geq 8$]{$2$-reflective lattices of signature $(n,2)$ with $n\geq 8$}

\author{Haowu Wang}

\address{Center for Geometry and Physics, Institute for Basic Science (IBS), Pohang 37673, Korea}

\email{haowu.wangmath@gmail.com}

\subjclass[2020]{11F55, 51F15, 17B67, 14J28}

\date{\today}

\keywords{Orthogonal modular forms, Reflection groups, Borcherds products, Reflective lattices}

\begin{abstract}
An even lattice $M$ of signature $(n,2)$ is called $2$-reflective if there is a non-constant modular form for the orthogonal group of $M$ which vanishes only on quadratic divisors orthogonal to $2$-roots of $M$. In 2017 Ma \cite{Ma17} proved that there are only finitely many $2$-reflective lattices of signature $(n,2)$ with $n\geq 7$. In this paper we extend the finiteness result of Ma to $n\geq 5$ and show that there are exactly forty-two $2$-reflective lattices of signature $(n,2)$ with $n\geq 8$.
\end{abstract}

\maketitle

\section{Introduction}
Let $M$ be an even lattice of signature $(n,2)$ with $n\geq 3$.  The type IV Hermitian symmetric domain $\cD(M)$ attached to $M$ is a connected component of the space
\begin{equation*}
\{[\mathcal{Z}] \in  \PP(M\otimes \CC):  (\mathcal{Z}, \mathcal{Z})=0, (\mathcal{Z},\bar{\mathcal{Z}}) < 0\}.
\end{equation*}
We denote by $\Orth^+ (M)$ the orthogonal group preserving $\cD(M)$ and $M$. Let $\Gamma$ be a finite-index subgroup of $\Orth^+(M)$ and $k$ be an integer. A holomorphic function $F$ on the affine cone 
$$
\cA(M)=\{ \mathcal{Z} \in M\otimes\CC : [\mathcal{Z}] \in \cD(M) \}
$$
is called a \textit{modular form} of weight $k$ and character $\chi$ for $\Gamma$
if it satisfies
\begin{align*}
F(t\mathcal{Z})&=t^{-k}F(\mathcal{Z}), \quad \forall t \in \CC^\times,\\
F(g\mathcal{Z})&=\chi(g)F(\mathcal{Z}), \quad \forall g\in \Gamma.
\end{align*}
A non-constant modular form $F$ is called \textit{reflective} if it vanishes only on quadratic divisors  
$$
l^\perp = \{ [\mathcal{Z}] \in \cD(M) : (\mathcal{Z}, l)=0 \}
$$
for some roots $l\in M$, that is, $l$ are primitive positive-norm vectors of $M$ whose associated reflection 
$$
\sigma_l: x\mapsto x-\frac{2(l,x)}{(l,l)}l, \quad x\in M
$$
fixes the lattice $M$, i.e. $\sigma_l \in \Orth^+(M)$. Bruinier's result \cite{Bru02, Bru14} yields that reflective modular forms can usually be constructed as automorphic Borcherds products \cite{Bor95, Bor98}.

Reflective modular forms first appeared in the works of Borcherds \cite{Bor95, Bor98} and Gritsenko--Nikulin \cite{GN98a, GN98b}. They have many important applications to generalized Kac--Moody algebras \cite{Bor92, Bor98, GN98a, GN98b, GN02, Sch06}, hyperbolic reflection groups \cite{Bor00, GN98a}, birational geometry of moduli spaces \cite{Bor96, BKP98, GN96, GN00, GHS07, GH14, Gri18, Ma18} and the classification and construction of free algebras of modular forms \cite{Wan21}. It is a common belief that reflective modular forms are very rare. In 1998 Gritsenko and Nikulin \cite[Conjecture 2.2.1]{GN98a} proposed the arithmetic mirror symmetry conjecture, stating that the number of lattices with a reflective modular form is finite up to scaling. Since then, many classifications of reflective modular forms have been obtained \cite{GN02, Bar03, Sch06, Sch17, Ma17, Ma18, Dit18, Wan18, GN18, Wan19, Wan22}. 

In this paper we study $2$-reflective modular forms, the most basic class of reflective modular forms. A reflective modular form on $\Gamma < \Orth^+(M)$ is called \textit{$2$-reflective} if its zero divisor is a linear combination of quadratic divisors $l^\perp$ for $l\in M$ with $(l,l)=2$. An even lattice $M$ is called \textit{$2$-reflective} if there is a $2$-reflective modular form for some finite-index subgroup of $\Orth^+(M)$. It follows from the symmetrization trick that if $M$ is $2$-reflective then there is a $2$-reflective modular form for $\Orth^+(M)$. Gritsenko and Nikulin observed \cite{GN96, GN00} that $2$-reflective modular forms are related to $K3$ surfaces and Calabi--Yau manifolds, in particular, they have a geometric interpretation as the automorphic discriminant of the moduli space of lattice-polarized $K3$ surfaces. 

There is some relation between $2$-reflective modular forms and hyperbolic $2$-reflection groups. Given a hyperbolic even lattice $S$. Let $W$ be the subgroup of $\Orth^+(S)$ generated by reflections associated with $2$-roots and $\mathcal{M}$ be an associated fundamental polyhedron. If the subgroup $A(\mathcal{M})$ of $\Orth^+(S)$ fixing $\mathcal{M}$ has finite index in the quotient group $\Orth^+(S) / W$ then $S$ is called \textit{$2$-reflective}. A  $2$-reflective hyperbolic lattice $S$ is called \textit{elliptic} if $A(\mathcal{M})$ is finite, otherwise it is called \textit{parabolic}. Nikulin and Vinberg \cite{Nik80b, Nik81, Nik84, Vin84, Nik96, Vin07} proved that the set of $2$-reflective hyperbolic lattices $S$ with $\rank(S)\geq 3$ is finite and gave a full classification of elliptic $2$-reflective hyperbolic lattices (see e.g. \cite[Section 3.2]{GN18} for a list). This classification was motivated by one result of Pjatecki\u{\i}-\v{S}apiro and \v{S}afarevi\v{c} \cite{PS71}, proving that a complex algebraic $K3$ surface with the Picard lattice $S$ has finite automorphism group if and only if $S$ is elliptic $2$-reflective. 

The arithmetic mirror symmetry conjecture of Gritsenko--Nikulin \cite[Section 2]{GN00} predicts that 
\begin{itemize}
    \item[(a)] there are only finitely many $2$-reflective lattices of signature $(n,2)$ with $n\geq 3$;
    \item[(b)] if $M$ has a $2$-reflective modular form $F$ then the hyperbolic lattice $c_M^\perp / \ZZ c$ is $2$-reflective for any primitive norm zero vector $c\in M$ such that $F$ vanishes on some $v^\perp$ with $v\in c_M^\perp / \ZZ c$.
\end{itemize}
Part (b) was proved in 2003 by Looijenga \cite[Corollary 5.11]{Loo03}. Part (a) was later proved in 2017 by Ma \cite{Ma17} for $n\geq 7$. Part (b) of the Gritsenko--Nikulin conjecture does not lead to an exact classification of $2$-reflective lattices, because the classification of parabolic $2$-reflective hyperbolic lattices is unknown. Ma’s proof is in algebraic geometry and his result is ineffective to classify $2$-reflective lattices. It turns out that one may need new ways to attack this problem.

In \cite{Wan19} the author developed an approach based on the theory of Jacobi forms \cite{EZ85, Gri18} to classify $2$-reflective lattices. Let $U$ be an even unimodular lattice of signature $(1,1)$ and $L$ be an even positive definite lattice. This approach yields that  if $2U\oplus L$ has a $2$-reflective modular form then either $L$ has no $2$-roots or $L$ contains a sublattice of the same rank generated by $2$-roots satisfying some strong constrains. After detailed analysis, it was found that there are exactly fifty-one $2$-reflective lattices of type $2U\oplus L$, where $L$ has $2$-roots. 

In this paper we improve Ma's result by means of Part (b) of the Gritsenko--Nikulin conjecture.

\begin{theorem}\label{MTH0}
There are only finitely many $2$-reflective lattices of signature $(n,2)$ with $n\geq 5$.     
\end{theorem}

We also give a complete classification of $2$-reflective lattices of signature $(n,2)$ with $n\geq 8$. 

\begin{theorem}\label{MTH}
There are exactly forty-two $2$-reflective lattices of signature $(n,2)$ with $n\geq 8$ up to isomorphism. They are formulated in Table \ref{tab:2-reflective}. In particular, there is no $2$-reflective lattice of signature $(n,2)$ for $n\geq 13$ and $n\neq 18, 19, 26$. 
\end{theorem}

Theorem \ref{MTH} has been proved in \cite[Theorem 1.1]{Wan19} for $n\geq 14$. In this paper we give a simpler proof and extend it to $n\geq 13$ (see Theorem \ref{th:n>=13}). To prove Theorem \ref{MTH} for $n\leq 12$, we combine the main results of \cite{Wan19} and some techniques of lattices so that we can drop the $2U$ assumption. The proof does not rely on the classification of $2$-reflective hyperbolic lattices. Note that there are $10$ lattices in Table \ref{tab:2-reflective} that do not appear in the table of \cite[Theorem 1.2]{Wan19}.

We remark that the converse of Part (b) of the Gritsenko--Nikulin conjecture does not hold. For example, there are indeed elliptic $2$-reflective hyperbolic lattices $S$ of rank $13\leq \rank(S) \leq 17$, but there is no $2$-reflective lattice of signature $(\rank(S), 2)$. We also remark that there are $2$-reflective lattices in Table \ref{tab:2-reflective} which induces parabolic $2$-reflective hyperbolic lattices through Part (b) of the Gritsenko--Nikulin conjecture, such as $U\oplus E_8(2)$, $U(2)\oplus 8A_1$ and $U\oplus E_6'(3)$. 

We have mentioned that elliptic $2$-reflective hyperbolic lattices are related to $K3$ surfaces with finite automorphism group. It would be interesting to know if $2$-reflective lattices in Table \ref{tab:2-reflective} correspond to a certain more special class of $K3$ surfaces (see e.g. \cite[Section 3]{GN00}). 

This paper is organized as follows. In Section \ref{sec:lemmas} we prove some technical lemmas about lattices and $2$-reflective modular forms. Section \ref{sec:Ma} is devoted to the proof of Theorem \ref{MTH0}. In Section \ref{sec:proof} we prove Theorem \ref{MTH} and give three corollaries.  

\begin{table}[ht]
\caption{$2$-reflective lattices of signature $(n,2)$ with $n\geq 8$}
\label{tab:2-reflective}
\renewcommand\arraystretch{1.1}
\[
\begin{array}{|l|l|}
\hline
n & \text{$2$-reflective lattice} \\
\hline 
26 & 2U\oplus 3E_8 \\
\hline 
19 & 2U\oplus 2E_8\oplus A_1 \\
\hline 
18 & 2U\oplus 2E_8 \\
\hline 
12 & 2U\oplus E_8\oplus 2A_1 \\
\hline 
11 & 2U\oplus D_4\oplus 5A_1, \quad 2U\oplus 2D_4\oplus A_1, \quad 2U\oplus D_8\oplus A_1, \quad 2U\oplus E_8\oplus A_1 \\
\hline
10 & 2U\oplus E_8, \quad 2U\oplus D_8, \quad 2U\oplus 2D_4, \quad 2U\oplus D_8'(2), \\
&2U\oplus E_7\oplus A_1, \quad 2U\oplus D_6\oplus 2A_1, \quad 2U\oplus D_4\oplus 4A_1, \quad 2U\oplus 8A_1, \\
&2U\oplus E_8(2), \quad U\oplus U(2)\oplus E_8(2), \quad U\oplus U(2)\oplus 8A_1, \quad 2U(2)\oplus 8A_1\\
\hline 
9 & 2U\oplus D_7, \quad 2U\oplus A_7, \quad 2U\oplus E_7, \quad 2U\oplus E_6\oplus A_1, \quad 2U\oplus D_6\oplus A_1, \\ 
&2U\oplus D_4\oplus 3A_1, \quad 2U\oplus 7A_1, \quad U\oplus U(2)\oplus 7A_1, \quad 2U(2)\oplus 7A_1 \\
\hline 
8 & 2U\oplus D_6, \quad 2U\oplus A_6, \quad 2U\oplus 2A_3, \quad 2U\oplus 3A_2, \quad 2U\oplus E_6, \\
&2U\oplus D_5\oplus A_1, \quad 2U\oplus A_5\oplus A_1, \quad 2U\oplus D_4\oplus 2A_1, \quad 2U\oplus 6A_1, \\
&2U\oplus E_6'(3), \quad U\oplus U(3)\oplus E_6'(3), \quad U\oplus U(2)\oplus 6A_1, \quad 2U(2)\oplus 6A_1 \\
\hline
\end{array} 
\]
\end{table}

\section{Basic lemmas}\label{sec:lemmas}
In this section we collect and prove some basic lemmas about lattices and $2$-reflective modular forms that we will use later. 

Let $M$ be an even lattice of rank $\rank(M)$ with a bilinear form $(-,-)$ and dual lattice $M'$. Let $A_M=M'/M$ denote the discriminant group of $M$. We denote the minimal number of generators of $A_M$ by $l(M)$ and the maximal order of elements of $A_M$ by $e(M)$. The integers $l(M)$ and $e(M)$ are called the length and exponent of $A_M$, respectively. Let us fix a basis of the (unique) even unimodular lattice of signature $(1,1)$ as 
$$
U=\ZZ e + \ZZ f, \quad (e,e)=(f,f)=0, \quad (e,f)=1.
$$
For any positive integer $a$, we denote by $M(a)$ the lattice with abelian group $M$ and rescaled bilinear form $a(-,-)$. The level of $M$ is the smallest positive integer $m$ such that $m(x,x)\in 2\ZZ$ for all $x\in M'$. An embedding $M_1 \hookrightarrow M$ of even lattices is called primitive if $M/M_1$ is a free $\ZZ$-module. A given embedding $M \hookrightarrow M_1$ of even lattices, for which $M_1/M$ is a finite abelian group, is called an even overlattice of $M$. For any $v\in M$ we define an ideal of $\ZZ$ as
$$
(v,M):=\{ (v,x) : x\in M \}.
$$
We use $A_n$, $D_n$, $E_6$, $E_7$ and $E_8$ to denote the usual irreducible root lattices (see \cite{CS99}). We refer to \cite{CS99} for the notion of the genus of a lattice. 

\begin{lemma}\label{lem:U}
Let $M$ be an even lattice of signature $(n,1)$ with $n\geq 2$. If the length of $A_M$ satisfies that $l(M)\leq n-2$, then there exists an even positive definite lattice $L$ such that $M=U\oplus L$.
\end{lemma}
\begin{proof}
It is a direct consequence of Nikulin's results \cite{Nik80} (see e.g. \cite[Lemma 2.3]{Wan19} for a proof).    
\end{proof}

\begin{lemma}\label{lem:maximal}
Let $M$ be a maximal even lattice of signature $(n,2)$ with $n\geq 5$. Then $M$ can be represented as $M=2U\oplus L$ for some even positive definite lattice $L$.    
\end{lemma}
\begin{proof}
Let $c$ be a primitive norm zero vector of $M$. Since $M$ is maximal, $(c,M)=\ZZ$, which yields a decomposition $M=U\oplus K$. Since $K$ has signature $(n-1,1)$ and $\rank(K)=n\geq 5$, there is a primitive norm zero vector of $K$ denoted $c_1$. Similarly, $(c_1,K)=\ZZ$ and $K=U\oplus L$ for some $L$. We then obtain the desired decomposition. 
\end{proof}

\begin{lemma}\label{lem:exponent}
Let $M$ be an even lattice of signature $(n,2)$ with $n\geq 8$. There exists an even overlattice $M_1$ of $M$ satisfying the following conditions
\begin{enumerate}
    \item $M_1$ can be represented as $2U\oplus L$;
    \item $A_M$ and $A_{M_1}$ have the same exponent, i.e. $e(M)=e(M_1)$; 
    \item the length of $A_{M_1}$ satisfies that $l(M_1)\leq 5$. 
\end{enumerate}
\end{lemma}
\begin{proof}
This follows from \cite[Lemma 1.7]{Ma18} and its proof.     
\end{proof}

\begin{lemma}\label{lem:2-roots}
Let $L$ be an even positive definite lattice of rank $\rank(L)$. If the $2$-component of $A_L$ has length $l(A_L)_2 \leq \rank(L)-3$ and the $p$-component of $A_L$ has length $l(A_L)_p\leq \rank(L)-2$ for any odd prime $p$, then there is a class in the genus of $L$ which has $2$-roots.
\end{lemma}
\begin{proof}
Recall that $U=\ZZ e + \ZZ f$ with $e^2=f^2=0$ and $(e,f)=1$.
We define $M=U\oplus L$. Let us fix $v=e+f$ and $u=e-f$. Note that $v^2=2$ and $u^2=-2$. The orthogonal complement of $v$ in $M$ has the form $M_v=\ZZ u \oplus L$, so it has signature $(\rank(L),1)$. By assumptions, we have
\begin{align*}
&l(A_{M_v})_2 = l(A_L)_2 + 1 \leq \rank(L)-2,\\
&l(A_{M_v})_p = l(A_L)_p \leq \rank(L)-2, \quad \text{for any odd prime $p$}.
\end{align*}
Therefore, $l(M_v)\leq \rank(L)-2$. 
By Lemma \ref{lem:U}, there exists an even positive definite lattice $L_0$ such that $M_v=U\oplus L_0$. Since $U\oplus L_0\oplus \ZZ v$ has an even overlattice isomorphic to $M$, there exists an even overlattice $T$ of $L_0\oplus \ZZ v$ satisfying $M\cong U\oplus T$. By construction, $v\in T$, so $T$ has $2$-roots. Thus $T$ gives a desired class in the genus of $L$.
\end{proof}

We recall some basic properties of $2$-elementary lattices. An even lattice $M$ is called \textit{$2$-elementary} if $A_M\cong (\ZZ / 2\ZZ)^a$ for some non-negative integer $a$. The genus of a $2$-elementary lattice is described by Nikulin \cite[Theorem 3.6.2]{Nik80}. In particular, we have the following.

\begin{lemma}\label{lem:2-elementary}
Let $M$ be a $2$-elementary even lattice of signature $(n,2)$ with $n\geq 3$. Suppose that $A_M \cong (\ZZ / 2\ZZ)^a$ for some non-negative integer $a$.  Then the following holds.
\begin{enumerate}
    \item $a\leq n+2$ and $n+a$ is even.
    \item There are at most two distinct $M$ up to isomorphism when $n$ and $a$ are fixed. 
    \item When $n, a$ are fixed and $4$ does not divide $n-2$, $M$ is unique up to isomorphism if it exists. 
\end{enumerate}
\end{lemma}

We now give some lemmas about $2$-reflective modular forms and $2$-reflective lattices.

\begin{lemma}[Lemma 2.3 in \cite{Ma17}]\label{Lem:reductionMa}
If $M$ is $2$-reflective, then any even overlattice of $M$ is also $2$-reflective.  If $M$ is not $2$-reflective, neither is any finite-index sublattice
of $M$.
\end{lemma} 

\begin{lemma}[Lemma 5.2 in \cite{Wan19}]\label{lem:pullback}
Let $M$ be an even lattice of signature $(n,2)$ with $n\geq 3$ and $L$ be an even positive definite lattice. If $M\oplus L$ is $2$-reflective, then $M$ is also $2$-reflective.    
\end{lemma}

We now introduce a particular class of $2$-reflective modular forms. A modular form for $\Orth^+(M)$ is called \textit{complete $2$-reflective} if its zero divisor is a linear combination of all quadratic divisors orthogonal to $2$-roots with multiplicity one. An even lattice is called \textit{complete $2$-reflective} if it has a complete $2$-reflective modular form. 

\begin{lemma}[Lemma 4.1 in \cite{Wan21}]\label{lem:complete}
Let $M=U\oplus U(m)\oplus L$. If $M$ is complete $2$-reflective then any even overlattice of $M$ is also complete $2$-reflective.    
\end{lemma}

\begin{lemma}\label{lem:non-complete}
Let $M=2U\oplus L$ be a $2$-reflective lattice. If $M$ is not complete $2$-reflective, then there exists an even lattice $K$ such that $M\cong A_1\oplus K$.   
\end{lemma}
\begin{proof}
By assumptions, there exists a $2$-root $v$ of $M$ with $(v,M)=2\ZZ$, because the set of $2$-roots $u\in M$ with $(u,M)=\ZZ$ is transitive under the action of $\Orth^+(M)$ (see \cite[Proposition 3.3]{GHS09}). We conclude from \cite[Lemma 7.5]{GHS13} that $M = \ZZ v \oplus M_v$, where $M_v$ is the orthogonal complement of $v$ in $M$.  We then prove the lemma.  
\end{proof}

\section{A proof of Theorem \ref{MTH0}}\label{sec:Ma}
Ma \cite{Ma17} proved that the set of $2$-reflective lattices of signature $(n,2)$ with $n\geq 7$ is finite. We improve Ma's result by a new method. 

\begin{theorem}
There are only finitely many $2$-reflective lattices of signature $(n,2)$ with $n\geq 5$.     
\end{theorem}
\begin{proof}
We first prove the theorem for $n\geq 7$, which reproves Ma's result.
Let $M$ be a $2$-reflective lattice of signature $(n,2)$ with $n\geq 7$. By \cite[Lemma 4.8]{Ma17} there exists an even overlattice $M_1$ of $M$ with length $l(M_1)\leq 4$ and exponent $e(M_1)=e(M)$ or $e(M)/2$. By Lemma \ref{lem:U} we can write $M_1=2U\oplus L$. Lemma \ref{Lem:reductionMa} yields that $M_1$ is $2$-reflective. Applying Part (b) of the Gritsenko--Nikulin conjecture (proved by Looijenga \cite{Loo03}) or Borcherds' result \cite[Theorem 12.1]{Bor98} to $2U\oplus L$, we find that $U\oplus L$ is a $2$-reflective hyperbolic lattice. Nikulin and Vinverg have proved that there are only finitely many $2$-reflective hyperbolic lattices. Therefore, both the exponents $e(M)$ and $e(M_1)$ are bounded from above. We then prove the desired result. 

We then consider the remaining cases. Let $M$ be a $2$-reflective lattice of signature $(n,2)$ with $n=5$ or $6$. According to \cite[Lemma 4.8]{Ma17}, there exists an even overlattice $M_1$ of $M$ such that $e(M_1)=e(M)$ or $e(M)/2$, $l(A_{M_1})_2\leq 4$ and $l(A_{M_1})_p \leq 3$ for any odd prime $p$. 

If there is a $2$-reflective modular form on $\Orth^+(M_1)$ with simple zeros, then we conclude from \cite[Corollary 1.10]{Ma18} that the number of such $M_1$ is finite up to isomorphism.  Therefore, the exponent $e(M)$ is bounded from above. We then prove the finiteness of $M$. 

Suppose that there is no $2$-reflective modular form on $\Orth^+(M_1)$ with simple zeros. We claim that $M_1$ has a $2$-reflective modular form $F$ which vanishes on some quadratic divisor $v^\perp$, where $v\in M_1$ with $(v,v)=2$ and $(v,M_1)=2 \ZZ$. Otherwise, there would be a modular form on $\Orth^+(M_1)$ whose zero divisor is a linear combination of quadratic divisors $l^\perp$ with some fixed multiplicity $m$, where $l$ takes over $2$-roots of $M_1$ with $(l,M_1)=\ZZ$, because the set of these $l$ is transitive under $\Orth^+(M_1)$. Since $M_1$ splits $U$, by \cite[Corollary 1.3]{Bru14} the modular form $F$ can be constructed as a Borcherds product on some sublattice of $M_1$. Therefore, there exists a modular form $F_1$ with simple zeros such that $F=F_1^m$. This contradicts the assumption. 

The existence of $v$ yields a decomposition $M_1=A_1\oplus K$ for some $K$ with $l(A_K)_p\leq 3$ for any prime $p$. Therefore, we can write $K=U\oplus T$ and thus $M_1=U\oplus T\oplus A_1$. By Part (b) of the Gritsenko--Nikulin conjecture, the hyperbolic lattice $T\oplus A_1$ is $2$-reflective. This implies the finiteness of $M_1$. We then finish the proof.
\end{proof}

\section{A proof of Theorem \ref{MTH}}\label{sec:proof}
In this section we present a proof of Theorem \ref{MTH}. The proof is divided into six cases. 

\begin{theorem}\label{th:n>=13}
The lattices $2U\oplus 3E_8$, $2U\oplus 2E_8\oplus A_1$ and $2U\oplus 2E_8$ are the only $2$-reflective lattices of signature $(n,2)$ with $n\geq 13$.     
\end{theorem}
\begin{proof}
It was proved by Ma \cite[Proposition 3.1]{Ma17} that $2U\oplus 3E_8$ is the unique $2$-reflective lattice of signature $(n,2)$ with $n\geq 26$. 
We now assume that $13\leq n \leq 25$.  

Suppose that $M$ is a maximal even lattice of signature $(n,2)$ and it is $2$-reflective. The length of $A_M$ satisfies that $l(M)\leq 3$. By Nikulin's results \cite[Corollaries 1.10.2 and 1.13.3]{Nik80}, we can write $M=E_8\oplus K$ for some maximal even lattice $K$. By Lemma \ref{lem:maximal}, we can further write $K=2U\oplus L$. Thus we have a decomposition $M=2U\oplus E_8 \oplus L$ with $3\leq \rank(L)\leq 15$.

By \cite[Theorem 6.2]{Wan19}, the sublattice $R$ of $E_8\oplus L$ generated by $2$-roots has the full rank $n-2$. Moreover, we can decompose $R$ into irreducible root lattices of type $ADE$ as
$$
R=E_8 \oplus R_1 \oplus mA_1,
$$
where $m$ is some non-negative integer and $R_1$ is a direct sum of some irreducible root lattices not of type $A_1$ contained in $L$. All irreducible components of $R$ not of type $A_1$ are required to have the same Coxeter number. Therefore, if $R_1$ is not zero, then it has to be $E_8$, because $\rank(R_1)\leq 15$. By the last statement of \cite[Theorem 6.2 (c)]{Wan19}, we have the expression 
$$
E_8\oplus L=2E_8\oplus (n-18)A_1 \quad \text{or} \quad E_8\oplus (n-10)A_1.
$$

In the former case, the assumption that $M$ is maximal forces that $n-18\leq 3$.  When $n=18$, $M=2U\oplus 2E_8$. When $n=19$, $M=2U\oplus 2E_8\oplus A_1$. When $n=20$, by Lemma \ref{lem:2-elementary} we have
$$
M=2U\oplus 2E_8\oplus 2A_1 \cong 2U\oplus E_8\oplus D_{10}.
$$
The second model of $M$ contradicts \cite[Theorem 6.2 (b)]{Wan19}, because $E_8$ and $D_{10}$ have distinct Coxeter numbers. When $n=21$, it follows from Lemma \ref{lem:pullback} that $M= 2U\oplus 2E_8\oplus 3A_1$ is not $2$-reflective.

In the latter case, the assumption that $M$ is maximal forces that $n-10\leq 3$. When $n=13$, Lemma \ref{lem:2-elementary} yields
$$
M=2U\oplus E_8\oplus 3A_1 \cong 2U\oplus E_7 \oplus D_4,
$$
which contradicts \cite[Theorem 6.2 (b)]{Wan19}, because $E_7$ and $D_{4}$ have distinct Coxeter numbers. 

We now consider the general case. Let $M$ be a $2$-reflective lattice of signature $(n,2)$ with $13\leq n\leq 25$. It remains to show that $M$ has to be maximal. 

Suppose that $M$ is not maximal and $M_1$ is a maximal even overlattice of $M$. As a maximal $2$-reflective lattice, $M_1$ has to be $2U\oplus 2E_8\oplus A_1$ or $2U\oplus 2E_8$ by the discussions above. In particular, $n=19$ or $18$. For such $n$, we can adapt the above argument to show that $2U\oplus 2E_8\oplus A_1$ and $2U\oplus 2E_8$ are the only $2$-reflective lattices $M$ of signature $(n,2)$ and length $l(M)\leq 3$. 

We claim that the order of the group $M_1/M$ is not a prime, otherwise the order of $A_M$ would be $2p^2$ or $p^2$. Thus $l(M)\leq 3$, which forces that $M=M_1$, a contradiction. Therefore, there exists an even lattice $M_2$ such that $M< M_2 < M_1$ and $M_1/M_2$ is a nontrivial cyclic group. It follows that $l(M_2)\leq 3$ and thus $M_2=M_1$, a contradiction. We then finish the proof.  
\end{proof}

\begin{theorem}\label{th:n=12}
The lattice $2U\oplus E_8\oplus 2A_1$ is the unique $2$-reflective lattice of signature $(12,2)$.     
\end{theorem}
\begin{proof}
Let $M$ be a $2$-reflective lattice of signature $(12,2)$. By Lemma \ref{lem:exponent}, there exists an even overlattice $M_1=2U\oplus L$ of $M$ satisfying that $e(M)=e(M_1)$ and $l(M_1)\leq 5$.  By Lemma \ref{lem:2-roots}, there exists a class $T$ in the genus of $L$ which has $2$-roots. Since $M_1\cong 2U\oplus T$ is $2$-reflective and $T$ has $2$-roots, we conclude from \cite[Theorem 1.2]{Wan19} that $M_1\cong 2U\oplus E_8\oplus 2A_1$. Therefore, both $M$ and $M_1$ are $2$-elementary. Thus $M'/M\cong (\ZZ/2\ZZ)^a$ for some positive integer $a$. By Lemma \ref{lem:2-elementary}, $a\leq 14$ and it is an even integer. For each such $a$ there is a unique lattice $M$ up to isomorphism. To prove the theorem it suffices to show that none of the following lattices is $2$-reflective:
\begin{align*}
&2U(2)\oplus 10A_1 < U(2)\oplus U\oplus 10A_1 < 2U\oplus 10A_1< \\
<&2U\oplus D_4\oplus 6A_1 < 2U\oplus D_6\oplus 4A_1 < 2U\oplus D_8\oplus 2A_1.     
\end{align*}
This follows from \cite[Theorem 1.2]{Wan19} and Lemma \ref{Lem:reductionMa}. 
\end{proof}

\begin{theorem}\label{th:n=11}
There are exactly four $2$-reflective lattices of signature $(11,2)$:
$$
2U\oplus D_4\oplus 5A_1, \quad 2U\oplus 2D_4\oplus A_1, \quad 2U\oplus D_8\oplus A_1, \quad 2U\oplus E_8\oplus A_1. 
$$
\end{theorem}
\begin{proof}
The proof is similar to that of Theorem \ref{th:n=12}.
Let $M$ be a $2$-reflective lattice of signature $(11,2)$. By Lemma \ref{lem:exponent}, there exists an even overlattice $M_1$ of $M$ with $e(M)=e(M_1)$ and $l(M_1)\leq 5$. By a similar argument, we have a decomposition $M_1=2U\oplus L_1$ for some $L_1$ having $2$-roots, and then we show that $M_1$ is isomorphic to $2U\oplus E_8\oplus A_1$, or $2U\oplus D_8\oplus A_1$ or $2U\oplus 2D_4\oplus A_1$. Therefore, $M$ is $2$-elementary. We write $A_M \cong (\ZZ/2\ZZ)^a$. By Lemma \ref{lem:2-elementary}, $a\leq 13$ and it is an odd integer. For each such $a$ there is a unique lattice $M$ up to isomorphism. It remains to prove that none of the following lattices is $2$-reflective:
$$
2U(2)\oplus 9A_1 < U(2)\oplus U\oplus 9A_1 < 2U\oplus 9A_1 \cong 2U\oplus E_8(2)\oplus A_1.
$$
We derive from \cite[Theorem 6.2]{Wan19} that $2U\oplus E_8(2)\oplus A_1$ is not $2$-reflective, because $E_8(2)$ has no $2$-roots. We then finish the proof of the theorem. 
\end{proof}

\begin{theorem}\label{th:n=10}
There are exactly twelve $2$-reflective lattices of signature $(10,2)$: 
\begin{align*}
&2U\oplus E_8& &2U\oplus D_8& &2U\oplus 2D_4& &2U\oplus D_8'(2)& \\
&2U\oplus E_7\oplus A_1& &2U\oplus D_6\oplus 2A_1& &2U\oplus D_4\oplus 4A_1& &2U\oplus 8A_1& \\
&2U\oplus E_8(2)& &U\oplus U(2)\oplus E_8(2)& &U\oplus U(2)\oplus 8A_1& &2U(2)\oplus 8A_1.&
\end{align*}
\end{theorem}
\begin{proof}
Let $M$ be a $2$-reflective lattice of signature $(10,2)$. By Lemmas \ref{lem:exponent} and \ref{lem:2-roots}, there exists an even overlattice $M_1=2U\oplus L$ of $M$ satisfying that $e(M)=e(M_1)$, $l(M_1)\leq 5$ and $L$ has $2$-roots.  By \cite[Theorem 1.2]{Wan19}, we find that $M_1$ is isomorphic to $2U\oplus E_8$, or $2U\oplus D_8$ or $2U\oplus 2D_4$, or $2U\oplus E_7\oplus A_1$, or $2U\oplus D_6\oplus 2A_1$. This implies that both $M$ and $M_1$ are $2$-elementary. We write $A_M\cong (\ZZ/2\ZZ)^a$. By Lemma \ref{lem:2-elementary}, $a\leq 12$ and it is an even integer. When $a=0$, $M=2U\oplus E_8$. For any even $a\geq 2$ there are exactly two lattices $M$ up to isomorphism: one with level $2$ and the other with level $4$. Since $2U(2)\oplus E_8(2)$ has no $2$-roots, it is not $2$-reflective. 
\end{proof}

The (unique) $2$-reflective modular form on $U\oplus U(2)\oplus E_8(2)$ was first constructed by Borcherds \cite{Bor96} in the study of the moduli space of Enriques surfaces. Borcherds also showed that this form defines the denominator of the fake monster Lie superalgebra (see \cite{Bor92}).  The $2$-reflective modular forms on lattices $2U(2)\oplus mA_1$ for $1\leq m\leq 8$ were constructed by Gritsenko--Nikulin \cite[Section 6.2]{GN18}. These forms are identical to some reflective modular forms of weight $12-m$ on $2U\oplus D_m$. 

The last two cases (i.e. $n=8, 9$) are more subtle because there are $2$-reflective lattices which are not $2$-elementary and we cannot use Lemma \ref{lem:exponent} in a direct way. 

\begin{theorem}\label{th:n=9}
There are exactly nine $2$-reflective lattices of signature $(9,2)$:  
\begin{align*}
&2U\oplus D_7& &2U\oplus A_7& &2U\oplus E_7& &2U\oplus E_6\oplus A_1& 
&2U\oplus D_6\oplus A_1& \\ &2U\oplus D_4\oplus 3A_1& &2U\oplus 7A_1& &U\oplus U(2)\oplus 7A_1& 
&2U(2)\oplus 7A_1.&
\end{align*}
\end{theorem}
\begin{proof}
Let $M$ be a $2$-reflective lattice of signature $(9,2)$. We fix a maximal even overlattice $M_0$ of $M$. Combining Lemmas \ref{lem:maximal} and \ref{lem:2-roots}, we have a decomposition $M_0=2U\oplus L_0$ such that $L_0$ has $2$-roots. Since $M_0=2U\oplus L_0$ is $2$-reflective and $L_0$ has $2$-roots, we conclude from \cite[Theorem 1.2]{Wan19} that $M_0$ is isomorphic to $2U\oplus E_6\oplus A_1$, or $2U\oplus E_7$ or $2U\oplus D_7$. Notice that $M<M_0<M_0'<M'$. There exist positive integers $t$ and $a_j$ for $1\leq j\leq t$ such that
$$
M' / M_0' \cong (\ZZ / a_1\ZZ) \oplus \cdots \oplus (\ZZ/a_t\ZZ).
$$
For any $a_s$ there exists an even overlattice $M_1$ of $M$ such that $M<M_1<M_0<M_0'<M_1'<M'$ and $M_1'/M_0' \cong \ZZ / a_s\ZZ$ (and thus $M_0/M_1\cong \ZZ / a_s\ZZ$).  We next discuss by cases.

\vspace{2mm}

\textbf{(I)} $M_0=2U\oplus E_6\oplus A_1$. We claim that $M=M_0$. 

Suppose that there are some $a_s>1$. Then $\det(M_1)=6a_s^2$ and $l(M_1)\leq 3$. By Lemma \ref{lem:2-roots}, there exists an even positive definite lattice $L_1$ with $2$-roots such that $M_1=2U\oplus L_1$. Thus $M_1$ lies in the table of \cite[Theorem 1.2]{Wan19} as a $2$-reflective lattice, which leads to a contradiction by comparing determinants of lattices. Therefore, every $a_j$ is $1$ and then $M= M_0= 2U\oplus E_6\oplus A_1$. 

\vspace{2mm}

\textbf{(II)} $M_0=2U\oplus D_7$. We claim that $M=M_0$. 

Suppose that there are some $a_s>1$. Then $M_1$ has determinant $4a_s^2$, length $l(M_1)\leq 3$ and exponent $e(M_1)\geq 4$. Similarly to the previous case, $M_1$ is a $2$-reflective lattice in the table of \cite[Theorem 1.2]{Wan19}, which leads to a contradiction by comparing determinants and exponents of lattices. 

\vspace{2mm}

\textbf{(III)} $M_0=2U\oplus E_7$. We claim that either $M=2U\oplus A_7$ or $M$ is $2$-elementary. 

A similar argument shows that every $a_j$ is either $1$ or $2$. Therefore, there exists a non-negative integer $a$ such that
$$
M'/M_0' \cong (\ZZ/2\ZZ)^a.
$$
A subgroup $G$ of $M'/M_0'$ of order $d$ corresponds to an even lattice $M_G$ of determinant $2d^2$ satisfying that $M<M_G<M_0$ and $M_0 / M_G \cong G$. More precisely, 
$$
M_G=\{ x\in M_0 : (x,y)\in \ZZ, \; y\in G+M_0'\}. 
$$

\textbf{(1)} When $a=1$, $\det(M)=2^3$, $l(M)\leq 3$ and thus we can write $M=2U\oplus L$ such that $L$ has $2$-roots. By \cite[Theorem 1.2 (c)]{Wan19}, $M$ is isomorphic to $2U\oplus A_7$ or $2U\oplus D_6\oplus A_1$. 

\textbf{(2)} We now consider the case $a\geq 2$. Let $G=\ZZ/2\ZZ \times \ZZ/2\ZZ$ be a subgroup of $M'/M_0'$. Similarly to the case $a=1$, we find that the lattice $M_1$ corresponding to a subgroup $\ZZ/2\ZZ$ of $G$ is $2U\oplus A_7$ or $2U\oplus D_6\oplus A_1$. Suppose that $M_1=2U\oplus A_7$. Then we have that $M<M_G<M_1$, $\det(M_G)=2^5$ and $l(M_G)\leq 3$. It follows that the $2$-reflective lattice $M_G$ has a decomposition $2U\oplus L_G$ such that $L_G$ has $2$-roots, which yields that $M_G$ lies in the table of \cite[Theorem 1.2 (c)]{Wan19}. This leads to a contradiction by considering the determinant and the length. Therefore, $M_1=2U\oplus D_6\oplus A_1$. We see from \cite[Theorem 6.2 (c)]{Wan19} that $M_1$ is not complete $2$-reflective, that is, every $2$-reflective modular form on $M_1$ either has a quadratic divisor with multiplicity larger than $1$ or does not vanish on some quadratic divisor orthogonal to a $2$-root of $M_1$.  

By Lemma \ref{lem:exponent}, there exists an even overlattice $M_2=2U\oplus L_2$ of $M$ satisfying that $e(M_2)=e(M)$ and $l(M_2)\leq 5$. We choose the above $M_0$ as a maximal even overlattice of $M_2$. 

If $l(M_2) \neq 1$, i.e. $M_2\neq 2U\oplus E_7$, then we can choose $M_1$ such that $M_2< M_1=2U\oplus D_6\oplus A_1$. By Lemma \ref{lem:complete}, the $2$-reflective lattice $M_2$ is not complete $2$-reflective. According to Lemma \ref{lem:non-complete}, we can write $M_2=A_1\oplus K$. Since $\det(M)=2^{2a+1}$, we have $l(M_2)=l(A_1)+l(K)$, so $l(K)\leq 4$. Therefore, by Lemma \ref{lem:U} we can write $K=2U\oplus T$. Since $M_2=2U\oplus T\oplus A_1$ is $2$-reflective, it lies in the table of \cite[Theorem 1.2 (c)]{Wan19}. We then conclude that both $M$ and $M_2$ are $2$-elementary. 

We complete the proof by the classification of $2$-elementary lattices. 
\end{proof}

\begin{theorem}
There are exactly thirteen $2$-reflective lattices of signature $(8,2)$:  
\begin{align*}
&2U\oplus D_6& &2U\oplus A_6& &2U\oplus 2A_3& &2U\oplus 3A_2& &2U\oplus E_6& \\
&2U\oplus D_5\oplus A_1& &2U\oplus A_5\oplus A_1& &2U\oplus D_4\oplus 2A_1& &2U\oplus 6A_1& &2U\oplus E_6'(3)& \\
&U\oplus U(3)\oplus E_6'(3)& &U\oplus U(2)\oplus 6A_1& &2U(2)\oplus 6A_1&
\end{align*}
\end{theorem}
\begin{proof}
Let $M$ be a $2$-reflective lattice of signature $(8,2)$. We fix $M_0$ as a maximal even overlattice of $M$. Since $l(M_0)\leq 3$, we can represent $M_0=2U\oplus L_0$. By Lemma \ref{lem:2-roots}, we can assume that $L_0$ has $2$-roots. Since $M_0=2U\oplus L_0$ is $2$-reflective and $L_0$ has $2$-roots, we know from \cite[Theorem 1.2 (c)]{Wan19} that $M_0$ is isomorphic to $2U\oplus D_6$, or $2U\oplus A_6$, or $2U\oplus E_6$, or $2U\oplus D_5\oplus A_1$. Note that $M<M_0<M_0'<M'$. There exist positive integers $t$ and $a_j$ for $1\leq j\leq t$ such that
$$
M' / M_0' = (\ZZ / a_1\ZZ) \oplus \cdots \oplus (\ZZ/a_t\ZZ).
$$
For any $a_s$ there exists an even lattice $M_1$ such that $M<M_1<M_0$ and $M_1'/M_0' \cong \ZZ / a_s\ZZ$.  We next discuss by cases.    

\vspace{2mm}

\textbf{(I)} $M_0 = 2U\oplus A_6$. We claim that $M=M_0$. 

The above $M_1$ has determinant $7a_s^2$ and length $l(M_1)\leq 3$. By Lemmas \ref{lem:U} and \ref{lem:2-roots}, we have a decomposition $M_1=2U\oplus L_1$ such that $L_1$ has $2$-roots. Therefore, the $2$-reflective lattice $M_1$ lies in the table of \cite[Theorem 1.2 (c)]{Wan19}. We then find that $a_s$ has to be $1$. 

\vspace{2mm}

\textbf{(II)} $M_0 = 2U\oplus D_5\oplus A_1$. We claim that $M=M_0$. 

The above $M_1$ has determinant $2^3a_s^2$ and length $l(M_1)\leq 4$. We notice that $M_0$ is not complete $2$-reflective (see \cite[Theorem 6.2 (c)]{Wan19}). By Lemma \ref{lem:U}, $M_1$ splits $2U$. Thus Lemma \ref{lem:complete} yields that $M_1$ is not complete $2$-reflective. It follows from Lemma \ref{lem:non-complete} that we has a decomposition $M_1=A_1\oplus K$ with $l(K)\leq 4$. Therefore, we can write $K=2U\oplus T$ and then $M_1=2U\oplus A_1\oplus T$ by Lemma \ref{lem:U}. Thus the $2$-reflective lattice $M_1$ lies in the table of \cite[Theorem 1.2 (c)]{Wan19}. We then see that $a_s=1$. 

\vspace{2mm}

\textbf{(III)} $M_0=2U\oplus E_6$. We claim that $M=2U\oplus A_5\oplus A_1$ or $M$ has level $3$. 

\textbf{(1)} Suppose that there are some $a_s=2$. We show that $M=2U\oplus A_5\oplus A_1$. 

A subgroup $\ZZ/2\ZZ$ of $M'/M_0'$ induces an even lattice $M_1$ with $\det(M_1)=12$ and $l(M_1)\leq 2$. Therefore, by Lemmas \ref{lem:U} and \ref{lem:2-roots} the $2$-reflective lattice $M_1$ has an expression $M_1=2U\oplus L_1$ such that $L_1$ has $2$-roots. \cite[Theorem 1.2 (c)]{Wan19} then yields that $M_1= 2U\oplus A_5\oplus A_1$. If $M\neq M_1$ then there exists an even lattice $M_2$ satisfying that $M<M_2<M_1$ and $l(M_2)\leq 4$. Since $M_1$ is not complete $2$-reflective, by Lemma \ref{lem:complete} $M_2$ is not complete $2$-reflective, so we can write $M_2=A_1\oplus K$ with $l(K)\leq 4$ by Lemma \ref{lem:non-complete}. Therefore, we can represent $M_2=2U\oplus A_1\oplus T$ by Lemma \ref{lem:U}. By \cite[Theorem 1.2 (c)]{Wan19}, such a $2$-reflective lattice $M_2$ does not exist, leading to a contradiction. Therefore, $M=M_1=2U\oplus A_5\oplus A_1$. 

\textbf{(2)} Suppose that there is no $a_j = 2$. If there is $a_s>3$, then $\ZZ/a_s\ZZ$ induces a lattice $M_1$ with $\det(M_1)=3a_s^2$ and $l(M_1)\leq 3$. Therefore, we can write $M_1=2U\oplus L_1$ such that $L_1$ has $2$-roots. Clearly, such $2$-reflective lattice $M_1$ does not exist by \cite[Theorem 1.2 (c)]{Wan19}, a contradiction. Thus we can assume that
$$
M'/M_0' \cong (\ZZ / 3\ZZ)^t. 
$$
We next show that $M$ has level $3$. 

We denote the generators of $M'/M_0'$ by $v_i$ for $1\leq i\leq t$. 
Any subgroup $\latt{v_i}\cong \ZZ/3\ZZ$ induces an even lattice 
$$
M_i=\{ x\in M_0: (x,v_i) \in \ZZ \}
$$
with $\det(M_i)=3^3$ and $l(M_i)\leq 3$. 
Note that $M_i'$ is generated by $M_0'$ and $v_i$. By Lemmas \ref{lem:U} and \ref{lem:2-roots}, we can express $M_i=2U\oplus L_i$ such that $L_i$ has $2$-roots, and therefore $M_i$ lies in the table of \cite[Theorem 1.2 (c)]{Wan19}. We find that $M_i \cong 2U\oplus 3A_2$, so $3(v_i,v_i)\in 2\ZZ$ and $3v_i \in M_0$.

When $t>1$, for $i\neq j$ we define an even lattice
$$
M_{ij}=\{ x\in M_0: (x,v_i)\in \ZZ, \; (x,v_j)\in \ZZ\}
$$
with $\det(M_{ij})=3^5$.
Note that the dual lattice $M_{ij}'$ is generated by $M_0'$, $v_i$ and $v_j$. 

If $M_{ij}'/M_{ij}$ has elements of order $9$, then $l(M_{ij})\leq 4$. By Lemma \ref{lem:2-roots}, we can write $M_{ij}=2U\oplus L_{ij}$ for some $L_{ij}$ with $2$-roots. \cite[Theorem 1.2 (c)]{Wan19} implies that such a $2$-reflective lattice $M_{ij}$ does not exist. Therefore, each non-zero element of $M_{ij}'/M_{ij}$ has order $3$. 

We have thus proved that $M_{ij}'/M_{ij}=(\ZZ/3\ZZ)^5$, which implies that $M_{ij}$ has level $3$ and thus $M_{ij}\cong 2U\oplus E_6'(3)$. Thus $3(v_i,v_j)\in \ZZ$. It is easy to verify by definition that $M$ is of level $3$. 

Thus $M=U\oplus U(3)\oplus E_6'(3)$, $2U\oplus E_6'(3)$, $2U\oplus 3A_2$ or $2U\oplus E_6$. The lattice $2U(3)\oplus E_6'(3)$ has no $2$-roots, so it is not $2$-reflective. We remark that the complete $2$-reflective modular form on $U\oplus U(3)\oplus E_6'(3)$ is identical to the $6$-reflective modular form on $2U\oplus 3A_2$ by \cite[Lemma 2.2]{Wan22}.
 
\vspace{2mm}

\textbf{(IV)} $M_0=2U\oplus D_6$. We claim that $M$ is $2$-elementary or $M=2U\oplus 2A_3$. 

We can write
$$
M'/M_0' \cong (\ZZ / 2^{a_1}\ZZ)^{b_1} \oplus \cdots \oplus (\ZZ / 2^{a_t}\ZZ)^{b_t},
$$
otherwise there is an even lattice $M_1$ satisfying that $M<M_1<M_0$, $\det(M_1)=2^2 a^2$ for some odd integer $a$ and $l(M_1)\leq 2$. Thus we can write $M_1=2U\oplus L_1$ such that $L_1$ has $2$-roots. The $2$-reflective lattice $M_1$ contradicts \cite[Theorem 1.2 (c)]{Wan19}.

Assume that $M\neq M_0$. Let $v\in M'$ with $2v\in M_0'$ and $v\not\in M_0'$. We define 
$$
M_1=\{ x\in M_0: (x,v)\in \ZZ \}.
$$
Then $M_1'$ is generated by $M_0'$ and $v$. Note that $\det(M_1)=2^4$. We discuss by three cases.
\begin{enumerate}
    \item $M_1'/M_1=(\ZZ/2\ZZ)^4$. We show that $M$ is $2$-elementary.
    
    As a $2$-elementary lattice, $M_1=2U\oplus D_4\oplus 2A_1$. By replacing $M$ with an even overlattice of the same exponent (see Lemma \ref{lem:exponent}), we can assume that $l(M)\leq 5$. Then $M$ splits $2U$. Since $M_1$ is not complete $2$-reflective, we know from Lemma \ref{lem:complete} that $M$ is not complete $2$-reflective. Combining Lemma \ref{lem:non-complete} and Lemma \ref{lem:U} we have a
    decomposition $M=A_1\oplus K$ with $l(K)\leq 4$ and thus a decomposition $M=2U\oplus A_1\oplus T$. We then determine $M$ by \cite[Theorem 1.2 (c)]{Wan19} and find that it is $2$-elementary.  
    \item $M_1'/M_1=(\ZZ/4\ZZ)\oplus (\ZZ/2\ZZ)^2$.  
    Then $l(M_1)=3$ and thus we can express $M_1=2U\oplus L_1$ such that $L_1$ has $2$-roots. There is no such $2$-reflective lattice by \cite[Theorem 1.2 (c)]{Wan19}. 
    
    \item $M_1'/M_1=(\ZZ/4\ZZ)^2$. We show that $M=M_1=2U\oplus 2A_3$.
    
    In this case, $l(M_1)=2$ and thus $M_1$ is a $2$-reflective lattice in the table of \cite[Theorem 1.2 (c)]{Wan19}. It follows that $M_1=2U\oplus 2A_3$. Assume that $M\neq M_1$. We take a lattice $M_2$ satisfying that $M<M_2<M_1<M_1'<M_2'<M$ and $M_2'/M_1'=\ZZ/2\ZZ$. When $l(M_2)\leq 3$, we can express the $2$-reflective lattice $M_2$ as $2U\oplus L_2$ such that $L_2$ has $2$-roots. By \cite[Theorem 1.2 (c)]{Wan19}, such $M_2$ does not exist. 
    
    Therefore, $l(M_2)>3$ and further $M_2'/M_2 \cong (\ZZ/4\ZZ)^2\oplus (\ZZ/2\ZZ)^2$. There are two cases:
    \begin{enumerate}
        \item $M_2 \cong U\oplus U(2)\oplus 2A_3$. We observe that $U\oplus U(2)\oplus 2A_3 \cong 2U\oplus L_2$ for some $L_2$ with $2$-roots. By \cite[Theorem 1.2 (c)]{Wan19}, such $M_2$ does not exist, a contradiction. 
        \item $M_2 \cong U\oplus A_1\oplus A_1(-1)\oplus 2A_3$. By Lemma \ref{lem:U}, we have $A_1(-1)\oplus 2A_3 \cong U\oplus T$ for some $T$. Therefore,
        $$
        U\oplus A_1\oplus A_1(-1)\oplus 2A_3 \cong 2U\oplus A_1\oplus T. 
        $$
        By \cite[Theorem 1.2 (c)]{Wan19}, such $M_2$ does not exist, a contradiction. 
    \end{enumerate}
\end{enumerate}
We finish the proof by the discussions above and the classification of $2$-elementary lattices. 
\end{proof}

At the end of this section, we give three corollaries of the main theorem. 

\begin{corollary}\label{cor1}
There are exactly $21$ complete $2$-reflective lattices of signature $(n,2)$ with $n\geq 8$ up to isomorphism. They are formulated as follows:
\begin{align*}
&2U\oplus 3E_8, \quad 2U\oplus 2E_8, \quad 2U\oplus E_8, \quad 2U\oplus E_8(2), \quad U\oplus U(2)\oplus E_8(2), \quad 2U\oplus D_8, \quad 2U\oplus 2D_4, \\  &2U\oplus D_8'(2), \quad 2U(2)\oplus 8A_1, \quad 2U\oplus D_7, \quad 2U\oplus A_7, \quad 2U\oplus E_7, \quad 2U(2)\oplus 7A_1, \quad 2U(2)\oplus 6A_1, \\  &2U\oplus D_6, \quad 2U\oplus A_6, \quad 2U\oplus 2A_3, \quad 2U\oplus 3A_2, \quad 2U\oplus E_6, \quad 2U\oplus E_6'(3), \quad U\oplus U(3)\oplus E_6'(3).
\end{align*}
\end{corollary}
\begin{proof}
It is a direct consequence of \cite[Theorem 6.9]{Wan19} and Theorem \ref{MTH}. The lattice $U\oplus U(2)\oplus mA_1$ is not complete $2$-reflective for $6\leq m \leq 8$, because its even overlattice $2U\oplus mA_1$ is not complete $2$-reflective (see Lemma \ref{lem:complete}). 
\end{proof}

The weights of complete $2$-reflective modular forms on $14$ of the above $21$ lattices are formulated in \cite[Table 2]{Wan19}. The complete $2$-reflective modular form has weight $12$ on $2U\oplus E_8(2)$ and $2U\oplus E_6'(3)$, weight $12-m$ on $2U(2)\oplus mA_1$ for $m=6,7,8$, weight $4$ on $U\oplus U(2)\oplus E_8(2)$ and weight $3$ on $U\oplus U(3)\oplus E_6'(3)$.

\begin{corollary}
Let $L$ be a primitive sublattice of the Leech lattice satisfying the $\mathrm{Norm}_2$ condition, that is, for any $\gamma \in L'/L$ there exists $v\in L+\gamma$ such that $(v,v)\leq 2$. 
If the rank of $L$ is greater than $5$, then $L$ is isomorphic to $E_8(2)$, $E_6'(3)$ or the Leech lattice. 
\end{corollary}
\begin{proof}
Let $\Lambda$ denote the Leech lattice.
By \cite[Section 5.1]{Wan19}, the pullback of the Borcherds form on $2U\oplus \Lambda$ defines a complete $2$-reflective modular form of weight $12$ on $2U\oplus L$.  Note that $L$ has no $2$-roots. The result then follows from the above corollary. 
\end{proof}

\begin{corollary}
Let $M$ be an even lattice of signature $(n,2)$ with $n\geq 8$. If the ring of integral-weight modular forms for the discriminant kernel
$$
\widetilde{\Orth}^+(M) = \{ g \in \Orth^+(M) : g(x) - x \in M, \; \text{for all $x\in M'$} \}
$$
is freely generated by $n+1$ forms, then $M=2U\oplus L$ for $L=E_8$, $D_8$, $D_7$, $A_7$, $E_7$, $D_6$, $A_6$ or $E_6$. 
\end{corollary}
\begin{proof}
Suppose the ring of modular forms for $\widetilde{\Orth}^+(M)$ is freely generated by forms $F_i$ of weights $k_i$ for $1\leq i\leq n+1$. By \cite[Theorem 3.5]{Wan21}, the Jacobian of these $F_i$ is a complete $2$-reflective modular form of weight $n+\sum_{i=1}^{n+1}k_i$.  We then complete the proof by Corollary \ref{cor1} and \cite[Theorem 4.4]{Wan21}. 
\end{proof} 

\bigskip

\noindent
\textbf{Acknowledgements} 
The author is supported by the Institute for Basic Science (IBS-R003-D1). The author thanks Shouhei Ma for valuable discussions.

\bibliographystyle{plainnat}
\bibliofont
\bibliography{refs}

\end{document}